OpenStudy (aravindg):

Find the integral:

OpenStudy (aravindg):

\[\large \int\limits_{0}^{1}\sqrt{x}e^{\sqrt{x}}\]

OpenStudy (aravindg):

@AccessDenied

OpenStudy (accessdenied):

what if you took sqrt(x) = x / sqrt(x) and u = sqrt(x)

OpenStudy (aravindg):

How does it help?

OpenStudy (accessdenied):

u = sqrt(x) in the exponential u^2 = x multiplied to the exponential du = 1/2 sqrt(x) dx , what we put into the denominator, to a constant multiple

OpenStudy (aravindg):

Which gives me 2u^3 e^u du Should I do integration by parts now?

OpenStudy (accessdenied):

Yea, integration by parts twice should get rid of your u^2 factor (should be u^2, because u = sqrt(x) and x = (sqrt(x))^2 = u^2), the e^u will stay and its integral is very easy to take in the end.

OpenStudy (aravindg):

I see. I was wondering if we needed to apply any properties of definite integral and solve it directly.

OpenStudy (accessdenied):

Hm.. I can't say I know of any specific neat properties that work with the original case. Finding the antiderivative and then substituting in the boundaries was my only idea here.

OpenStudy (aravindg):

Okay. Thanks!

OpenStudy (accessdenied):

I stumbled across something while looking if there was an easier way of going about it (I've run into the x^n e^x case often, but always had to integrate by parts) If you've never seen it, it was called a 'tabular method of repeated integration by parts' and looked useful for doing these annoying multi-step integration by parts. it is news to me at least, lol. http://math.ucsd.edu/~wgarner/math20b/int_by_parts.htm

OpenStudy (aravindg):

Yes that's new to me too! Thanks for sharing :)